OpenStudy (anonymous):

find the slope between the following points (0, 3) and (8, -4)

OpenStudy (anonymous):

The Slope is -7/8 :D Medal pls haha

OpenStudy (anonymous):

the formula os slope is y2-y1/x2-x1 y2=-4/ y1=3 and x2= 8, x1 = 0 use the formula y2-y1/x2-x1 -4-3/8-0 -7/8

OpenStudy (anonymous):

you dont have to use that formula because i found it more confusing all you have to do is look at the 2 points, then from left to right see how far you have to go down and how far you have to go over. So in this question I counted down 7 (thats why it is -7) and over 8...Hope it helps :D